A roller coaster starts from rest at the top of a track 30.0 m long and inclined at 20.0° to the horizontal. Assume that friction can be ignored. A) what is the acceleration of the car?
B)How much time elapses before it reaches the bottom of the track?
1
Expert's answer
2020-11-12T09:15:50-0500

Answer

length of track(s) =30 m

Inclination "\theta=20.0°"

Car will have acceleration due to gravity

"a=gsin\theta\\=9.8\times sin20°\\=9.8\times 0.34\\a=3.33m/s^2"


Time taken by car to reach the bottom can be given as

"t=\sqrt{\frac{2s}{a}}\\t=\sqrt{\frac{2\times30}{3.33}}\\t=18.01second\\t=18s"